### v2

# unPIC: A Geometric Multiview Prior for Image to 3D Synthesis

Abstract
We introduce a hierarchical probabilistic approach to go from a 2D image to multiview 3D: a diffusion “prior” predicts the unseen 3D geometry, which then conditions a diffusion “decoder” to generate novel views of the subject. We use a pointmap-based geometric representation to coordinate the generation of multiple target views simultaneously. We construct a predictable distribution of geometric features per target view to enable learnability across examples, and generalization to arbitrary inputs images. Our modular, geometry-driven approach to novel-view synthesis (called “unPIC”) beats competing baselines such as CAT3D, EscherNet, Free3D, and One-2-3-45 on held-out objects from ObjaverseXL, as well as unseen real-world objects from Google Scanned Objects, Amazon Berkeley Objects, and the Digital Twin Catalog.
1 Introduction
Recovering 3D appearance from a single image is a hard, underspecified problem [13, 29]. A process that generates multiple views from a monocular image of a subject can serve as a useful precursor for 3D reconstruction [31, 12, 27]. Such a multiview prior is commonly implemented as a diffusion denoising probabilistic model [17] to capture a distribution of valid 3D outputs. A common pitfall, however, is that most implementations conflate the prediction of the unseen geometry and texture, making the output space complex and the task harder.
[…]
Our hierarchical approach is analogous to the unCLIP approach behind DALL-E 2 [39], a popular model for text-to-image synthesis: they first map CLIP text embeddings to image embeddings using a probabilistic prior, then decode the image embeddings to pixels. Our work is not only inspired by unCLIP’s two-stage generative process, but can also be seen as downstream of unCLIP. We call our approach “unPIC” – undo-a-picture with Probabilistic Inverse Cameras – to evoke the inverse of photography.
To specify the target shape of an object before filling in appearance-level details, we use a pointmap-based intermediate representation. We adapt the idea of a Normalized Object Coordinate Space (NOCS [45]), a function that maps every point on an object’s surface to a unique RGB color based on its spatial coordinates. When rendered as images (aka pointmaps), NOCS establishes point-to-point correspondence across viewpoints (Fig. 1). It is naturally suited to help coordinate the generation of consistent novel-view images. We introduce a version of NOCS that exploits an available degree of freedom—the orientation of the color space—to make the pointmaps more predictable for the prior module. Our version, called “CROCS”, enables camera-pose-invariant predictions, which in turn enables in-the-wild predictions.
Figure 1: Top: A hierarchical approach to novel-view synthesis. A prior models multiview geometric features from a single image and are jointly decoded to the target novel-view images. Our intermediate features, CROCS, establish point-to-point correspondence across views. Bottom: Samples from the prior and decoder. Our model exhibits transferrable shape understanding having never seen a real-world pixel.
Ultimately, our goal is to produce realistic and diverse shape and appearance completions from a single 2D image. To motivate our design, we present the following experiments: we first show that a diffusion-based appearance decoder achieves better novel-view synthesis when equipped with ground-truth NOCS pointmaps (Sec. 4.1). Next, we train a diffusion-based geometric prior to predict multiview pointmaps from a given source image (Sec. 4.2). Our camera-relative version of NOCS significantly improves the diffusion prior’s performance, achieving a 4x boost in shape prediction accuracy. Finally, we stack the separately trained prior and decoder for hierarchical novel-view synthesis from a single image (Sec. 5). We perform favorably against multiple baselines CAT3D, One-2-3-45, OpenLRM, EscherNet, and Free3D on appearance and shape prediction metrics. These results demonstrate our method’s superior shape understanding, and the benefit of geometry-driven multiview synthesis.
2 Related Work
Novel view synthesis: Optimization-based NVS methods such as Neural Radiance Fields [32] or Gaussian Splatting [26] fit a geometric representation to a specific scene. Such methods need several consistent input images to be able to interpolate between camera poses. When only one image is available, we need a learned prior to predict what an object/scene looks like from alternative views. One line of work builds on spatially arranged implicit representations—such as convolutional feature grids in pixelNeRF [54] or attention-based triplane tokens in LRM [18]—to enable feed-forward inference. These original methods are deterministic, thus suffering from averaged/blurry reconstructions, as well as inconsistent 3D outputs from different views of the same object. Probabilistic extensions such as DMV3D [53] help instantiate multiple versions of an object from a single image. In a similar vein, methods such as DreamFusion [37], Score Jacobian Chaining [46], Zero-1-to-3 [31], and ZeroNVS [41] use diffusion-based priors to iterate between novel view synthesis and 3D reconstruction. Such methods tend to be slow due to scene-specific optimization (typically NeRF- or SDF-based). They also suffer from issues [2] due to bootstrapping from pretrained text-to-image models.
Recently, it has become common to generate multiple novel views before reconstructing 3D. EscherNet [27], Free3D [57], ReconFusion [50], and CAT3D [12] all take this approach, but avoid using or producing any geometric explanation during the multiview synthesis. Although generating multiple images jointly helps boost consistency on its own, the lack of geometric grounding/supervision makes these methods susceptible to producing plausible images but implausible geometry. And while diffusion-based methods like SODA’s latent bottleneck [21] and 3DiM’s stochastic conditioning [49] can help reduce inconsistencies among generated novel views, they lack explicit geometric control. Our approach allows direct geometric supervision, and ensures that generated novel views are synchronized via an explicit geometric input. At inference time, our method does not depend on having access to any geometric information. We show our method achieves better consistency among novel views relative to geometry-free approaches.
Geometry-conditioned generative models: ControlNet [55] popularized the use of pixel-aligned hints to condition and guide a diffusion model. Other applications of geometry conditioning include motion brush animation [34] and pose-conditioned multi-object generation [22, 51]. This approach is especially tempting as diffusion models can be trained with conditioning dropout to enable classifier-free guidance [16], i.e., model a conditional and unconditional distribution simultaneously. At inference time, the model can cope with not having any geometric information, and still sample from the unconditional or source image-conditioned distribution. Few prior works (with the prominent exception of Motion-I2V [43]) have explored sampling geometric features using a diffusion prior, and feeding them to a decoder in place of true features or user inputs. We show this is not only possible, but desirable. Sampling from a prior ensures the geometric conditioning is realistic for the given example, diverse (e.g., different 3D shapes corresponding to a 2D image), and obtainable without human effort.
Pointmaps: Point-based representations are popular in 3D modeling. Recently, DUSt3R [47] introduced a novel approach to 3D reconstruction, estimating dense pointmaps to unproject a collection of images to 3D. DUSt3R and our work share two similarities: First, DUSt3R estimates pointmaps for two images at once to establish their correspondence. Likewise, we generate 8 pointmaps simultaneously to ensure they correspond with each other. Second, to handle scale ambiguity, DUSt3R normalizes pointmaps using their norm. This is similar to the scale normalization performed by NOCS. Two key differences between DUSt3R and our work are as follows: (i) We use a diffusion-based probabilistic model to handle the underspecified nature of image-to-3D. (ii) In DUSt3R, pointmaps are initially expressed in pairwise reference frames. These need to be aligned globally via optimization. In contrast, we express all pointmaps in a shared reference frame (i.e., CROCS color cube) from the outset. We go one step further—we set the orientation of the color cube based on the camera pose of the source image in each example. This makes the coloring predictable from arbitrary viewpoints, and learnable across examples.
Diffusion-based geometric priors: Pretrained diffusion models been shown to yield latents that are useful for various vision tasks including depth estimation and semantic correspondence [10, 9]. They have also been successfully trained to output dense annotations such as depth and surface normals [14, 25, 11]. Typically these models annotate monocular images rather than predict novel views. They lack geometric consistency when applied to different views of a scene. In fact, there is also recent work (called SpaRP [52]) which attempts to annotate images with NOCS-based pointmaps. The main difference with our work is that they only estimate NOCS for a given set of input images to infer their correspondence, rather than use NOCS to model the unseen geometry at novel views.
3 Method
3.1 A Hierarchical Multiview Framework
unPIC is a hierarchical diffusion model that takes a 2D image and predicts a set of target novel-view images. Given a single view of an object, unPIC comprises: (i) A prior that infers representations of the subject’s 3D geometry. This inference step combines the task of annotating the source image as well as predicting representations for novel views. Together, these predictions form a loop in representation space. (ii) An appearance decoder that transforms all multiview geometry predictions into textured images of the subject.
Annotating the source image alone may be a deterministic step (e.g., if using features from a typical deterministic encoder). In contrast, predicting novel-view features is inherently non-deterministic, due to the partial information observed in a single image. There may be several valid loops that go through the same point that describes the source image (see Figure 2). Hence, a probabilistic prior is essential, as opposed to the bulk of prior work [54, 18, 23]. As for the appearance decoder—which takes predicted representations and maps them to images—that could be (i) either deterministic or stochastic; (ii) either separately conditioned on each single view or jointly on multiviews. Our proposed representation (CROCS) captures the geometry but leaves the texture of the subject entirely unknown. Hence we use a probabilistic model again.
Figure 2: Schrödinger’s cup: two sets of valid novel views, following different trajectories in representation space. The observed view does not reveal whether the cup has a handle or not.
Depending on the choice of representation, the target annotations may vary haphazardly as a function of the camera pose, or remain invariant (e.g., global semantic features such as the class). A suitable representation must ensure the loops are well structured, in that there is a learnable relationship between target views across examples. This increases the benefit of modeling multiple views simultaneously, as we will show empirically in Table 2.
3.2 Target-Camera Poses for Arbitrary Source Poses
Our aim is to generate novel views predicting an object’s appearance as it spins around its vertical axis. We want to do this for arbitrary source images. In keeping with prior work [12], we predict a small set of maximally informative target poses, thereby enabling downstream 3D reconstruction or camera interpolation.
Concretely, we predict novel views at K camera poses, relative to any given image. These correspond to rotations $\theta_{k}=2\pi(k-1)/K$ of the object about its vertical axis, which remains aligned with the vertical axis of the world. We include the source camera pose as one of the targets, $\theta_{1}=0$; this helps anchor the model’s predictions of novel views. One can also vary the camera elevation angle (as in [52]) along with the azimuthal angle, but to allow comparison with competing methods, we choose not to.
Note that using $K$ camera poses does not equate to a fixed multi-camera rig (as in [3]). Rather, we allow the source camera to vary freely on hemispheres of arbitrary size around the object. Given any source camera, our target cameras are at the same height as the source, the same distance to the object, but at deterministic rotations around the object’s vertical axis, and facing the object.
Since we define target camera poses relative to the source camera, we do not need to explicitly condition our model on camera poses. This precludes camera control (the ability to generate a novel view at an arbitrary pose).
3.3 CROCS: Camera-Relative Object Coordinates
We predict the geometry of an object before decoding its multiview appearance. To do so, we rely on NOCS, a scale-free representation of geometry. The object/scene is first uniformly scaled to fit a 3D unit cube, so all spatial coordinates are in $[0,1]$. These can therefore be interpreted as RGB colors and rendered as images, which maps naturally onto our multiview framework. While a single NOCS image serves as a dense image annotation, a set of NOCS images taken together forms a point cloud.
Unfortunately, NOCS on its own does not provide a predictable coloring of the object’s surface—any rotation of the RGB color space would still produce a valid coloring of a fixed object. We introduce a version of NOCS that is canonicalized to paint an object relative to the source camera pose (see Figure 3). Say the source camera moves through a rotation of $\theta$ about the object’s vertical axis. We correct for that by applying an SO2 rotation to the Red and Green channels—which span the NOCS ground plane—of all pointmaps for the given object. Note that this destroys any pose relations across objects (our training data does not, for instance, paint all cups’ handles with a consistent color).
We address two geometric subtleties in working with CROCS in Appendix B—the first one is about renormalizing pointmaps after a change of reference frame to ensure they are still bounded in $[0,1]$. The second one explains our choice not to canonicalize CROCS colors with respect to varying camera elevations $\phi$.
Our CROCS representation diverges from the typical NOCS formulation [45, 28] a second way: when dealing with a multi-object scene, we treat it as a single object. This is better suited for our task, as we care less about capturing the pose of individual objects, but more about the spatial arrangement and relative position of different objects.
Figure 3: Camera-Relative Object Coordinate Spaces. We show two data-points (Left and Right columns) obtained from one object. Top-left: The wireframe shows the RGB reference cube used to paint the object surface. The large camera denotes the source view, whereas the smaller cameras denote (3 of 7) novel views. Top-right: Say all camera locations are rotated by $\theta=120$ degrees around the vertical axis (the object stays fixed). Then we also rotate the color reference cube by the same degree. This ensures each camera faces the same side of the cube that it was facing prior to the cameras’ rotation. Bottom: Target CROCS images in clockwise order corresponding to the cameras above. In a given data-point, any part of the object is consistently colored across target images. Across data-points, although a given part of the object may change colors, its color is predictable based on its location(s) in the target image(s). Each target view has a consistent color bias that is learnable across examples.
3.4 Dataset
We train on a combination of object and scene assets from Objaverse [5] and the alignment-finetuning subset of ObjaverseXL [6] . We holdout 50k randomly sampled objects from ObjaverseXL for evaluation. For each asset, we pre-render images from a dense grid of camera poses (at various $\theta$, $\phi$, and $r$; see Section 3.2). We also pre-render one NOCS pointmap per image using a fixed NOCS RGB reference frame. At training time, we adapt the pre-rendered NOCS pointmaps to CROCS on demand based on which camera is chosen as the source, using rotations of the RGB space as mentioned in Section 3.3.
When rendering images, we add an elevated point light behind the camera. We ensure it stays fixed relative to the camera, rather than fixed in the scene. This ensures the model cannot use lighting cues; it treats each image the same in terms of lighting. (A small number of assets have preexisting lights. We do not remove these.).
3.5 Diffusion Model
We use a standard pixel-based Diffusion Denoising setup [17]. The prior and decoder models use an identical architecture, broadly following the UNet from [12]. We train the two models independently to maximize the geometric diversity of the prior—else it would be forced to output the most optimal representations for the decoder’s loss (e.g., hacks to copy bits of texture from the source image). Moreover, we train the models from scratch, without relying on pre-trained weights, to avoid ambiguity about where the models gain their representational and reconstruction abilities.
To denoise $K=8$ multiview images simultaneously, we find that tiling them to form a superimage works best. We tile clockwise along 4 columns and 2 rows—this ensures that adjacent tiles always correspond to adjacent camera poses in 3D (see  Figure 3). The source image is also provided to the model as a superimage, but with 7 of 8 tiles remaining empty. For $K=4$, the superimage grid size is 2x2.
We run preliminary experiments (Sections 4.1 and 4.2) at base image size 64x64, predicting either 1, 4, or 8 targets. We use a standard cosine diffusion schedule [33] to train these models. Our final model works on 256x256 base images, predicting 8 targets at a superimage resolution of 1024x512. To scale to this size, we use the interpolated, logSNR-offset schedule from Eq. 6 of [19]. The higher noise levels are necessary for diffusion to overcome the correlations at this resolution.
4 Preliminary Experiments
4.1 What features are most useful to decode novel views?
To motivate geometric representations and the use of CROCS specifically, we run an experiment to decode a single novel view of an object at a fixed rotation ($\theta=90$ degrees). We train diffusion models that take a source image and a ground-truth annotation of the target image to predict the target image. The annotations range from geometric (like depth maps and CROCS), image-space (alpha masks), through feature maps (from DINOv2 [35] and CLIP [38]). All annotations are treated as images—they are upscaled (if required) and concatenated to the input latent before it is fed to the UNet for denoising. We run these experiments at 64x64 image size to allow multiple seeds.
We find in Table 1 that geometric annotations consistently outperform other choices. Given a source image, there is more ambiguity about the target shape than its texture or appearance, explaining the empirical results. We also see a benefit of CROCS over NOCS: coloring the annotation consistently makes it more useful than feeding in arbitrary colors. The full benefit of CROCS will be revealed when we try to predict CROCS and NOCS each in Sec 4.2.
Multiple targets. To examine the benefit of geometric features for multiview prediction, we run experiments with and without ground-truth annotations when predicting 4 or 8 target views. The target superimage dimensions are 128x128 and 256x128, respectively. We carefully control the diffusion architecture, applying attention at two downsampled resolutions in each case (when the feature height dimension is 16 or 8).
In Table 2, we see a large improvement in the MSE by conditioning on multiview CROCS pointmaps. In fact, we see increasing benefits from $K=1$ (using CROCS is $3.9$x better) to $K=4$ ($4.5$x better) and $K=8$ ($5.6$x better). The CROCS-conditional error is not only smaller relative to the unconditional case, it actually decreases in absolute terms when predicting more views. This is remarkable because the number of tokens doubles with the number of views, making it harder to model spatial relationships across the whole superimage (and perhaps leading to the drop in the unconditional case). These results help establish the advantage of (ground-truth) CROCS annotations in coordinating multiview synthesis.
Table 1: Single-target diffusion decoder experiments. We predict a fixed 90-degree object rotation from a given source image, with the help of a ground-truth annotation of the target image. We report the pixel-space Mean Square Error on random source views across 10k heldout objects. Stds. are across 3 training runs.
Annotation Type MSE (1e-3)
Source image only 19.042 $\pm$ 0.722
+ CLIP ViT L/16 (24, 24, 1024) 6.059 $\pm$ 0.448
+ DINOv2 B/14 (16, 16, 768) 5.734 $\pm$ 0.815
+ Alpha mask 8.169 $\pm$ 0.392
+ Depth map (from NOCS) 5.580 $\pm$ 0.849
+ NOCS 5.318 $\pm$ 0.070
+ CROCS 4.914 $\pm$ 0.224
Table 2: Multiple-target diffusion decoder experiments using ground-truth annotations. We report the pixel MSE (1e-3) on $K-1$ novel views as in Table 1. Stds. are across 5 training runs.
Annotation Type K=4 views K=8 views
Source image only 4.739 $\pm$ 0.215 5.817 $\pm$ 0.230
+ CROCS (K views) 1.065 $\pm$ 0.064 1.039 $\pm$ 1.331
4.2 What multiview features are predictable from a single image?
Having used ground-truth representations in Section 4.1, we now move to estimating multiview geometric representations from a single source image. We train a diffusion prior to predict pointmaps at 8 target views. We compare two choices of representations for how predictable they are from arbitrary source views: (i) NOCS images in a static reference frame, and (ii) CROCS, where we change the reference frame based on the source camera.
We find a 4x increase in predictability from using CROCS (see Table 3). Without CROCS, choosing a random source camera leaves the colors of the target images unpredictable because we do not condition the model on any camera extrinsics. With CROCS, each target pose has a biased distribution of colors, consistent across objects. This makes the geometric feature space unambiguous, and each representation loop more predictable, in turn freeing up the model’s probabilistic capacity for the ambiguous task of predicting unseen object/scene geometry.
Table 3: Diffusion prior experiments. We predict pointmaps at 4 and 8 target views, and report the pointmap MSE (1e-3).
Annotation Type K=4 views K=8 views
NOCS 11.94 15.82
CROCS 1.21 3.92
Table 4: Comparison with baselines on 4 datasets.
Dataset Method PSNR $\uparrow$ IoU $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ CLIP MV $\downarrow$
Objaverse XL CAT3D 22 0.63 84 0.23 0.75 2.44
EscherNet 12 0.62 79 0.26 0.65 2.36
Free3D 15 0.66 82 0.22 0.71 3.10
One-2-3-45 12 0.58 68 0.25 0.66 2.55
OpenLRM 14 0.62 103 0.23 0.71 2.38
unPIC-latents (ablation) 23 0.71 78 0.18 0.79 2.23
unPIC-pixels (ours) 25 0.79 83 0.19 0.78 2.16
Google Scanned Objects CAT3D 24 0.68 78 0.23 0.75 3.08
EscherNet 14 0.73 72 0.25 0.67 2.79
Free3D 16 0.73 83 0.23 0.71 3.60
One-2-3-45 13 0.65 57 0.25 0.66 2.74
OpenLRM 14 0.67 101 0.25 0.69 2.93
unPIC-latents (ablation) 23 0.80 69 0.18 0.78 2.42
unPIC-pixels (ours) 26 0.89 87 0.18 0.79 2.63
Amazon Berkeley Objects CAT3D 22 0.64 70 0.26 0.71 2.37
EscherNet 11 0.65 70 0.30 0.60 2.24
Free3D 14 0.73 64 0.24 0.66 2.87
One-2-3-45 11 0.61 48 0.27 0.62 2.22
OpenLRM 12 0.59 115 0.29 0.65 2.55
unPIC-latents (ablation) 25 0.83 50 0.16 0.78 1.88
unPIC-pixels (ours) 27 0.87 58 0.17 0.78 1.74
Digital Twin Catalog CAT3D 23 0.74 68 0.20 0.76 2.70
EscherNet 14 0.75 71 0.22 0.69 2.29
Free3D 17 0.80 75 0.19 0.72 2.86
One-2-3-45 14 0.74 49 0.20 0.70 1.98
OpenLRM 14 0.71 88 0.22 0.72 2.40
unPIC-latents (ablation) 25 0.84 60 0.14 0.82 1.89
unPIC-pixels (ours) 27 0.91 68 0.16 0.81 1.81
5 Main Results
We compare unPIC on single-image-to-multiview synthesis with the following state-of-the-art methods:
  1. 1.
    CAT3D (512x512): a SoTA-quality, geometry-free diffusion model trained using masked modeling of views on diverse datasets. While the original model is closed source, we received a Colab and checkpoint from the authors. The model uses a coarse view-sampling strategy to generate 7 anchor views given 1 source image. Subsequently, it conditions on the anchor views to generate a finer set of views using the same model. We focus on the coarse stage for our evaluation.
  2. 2.
    EscherNet (256x256): a geometry-free diffusion model equipped with a specialized camera encoding (CaPE), allowing a flexible number of input images and target views.
  3. 3.
    Free3D (256x256): a geometry-free multiview diffusion model that uses Plücker rays to encode target camera poses and modulate the diffusion latents. It is trained on Objaverse only.
  4. 4.
    One-2-3-45 (256x256): a SoTA version of Zero-1-to-3 [31]. Its diffusion-based image prior takes camera rotation and translation (R and T) parameters to transform a given image. Each novel view is sampled individually, leaving the outputs prone to multiview inconsistencies. We focus on evaluating the geometry-free stage, feeding R and T parameters to match our target poses, but using the model’s own estimation of the camera height.
  5. 5.
    OpenLRM (288x288): a geometry-aware but deterministic method based on the original LRM [18]. It uses a Transformer to convert DINO features into implicit triplane representations of 3D shape. The triplane tokens are spatially indexed/queried to parameterize a NeRF model of the scene, mitigating view consistency issues and allowing image-generation at arbitrary camera poses.
Figure 4: Qualitative comparison. The left column shows the source image, while the remaining images in each row are predicted novel views. One-2-3-45 produces multiview inconsistencies. CAT3D can squash the shapes in unseen views.
Our model is based on pixel-diffusion and trained from scratch on synthetic data only. In contrast, most competing methods have the advantage of starting from a pretrained model: Free3D, EscherNet, and Zero-1-to-3 start from Stable Diffusion; CAT3D starts from a similar pre-trained latent diffusion model [40]; and OpenLRM bootstraps from DINO features. The large-scale pretraining of these models/features is an obvious advantage.
An ablation. To demonstrate that training from scratch on synthetic data is not necessarily a disadvantage, we run the following baseline: we train a latent diffusion model (denoted unPIC-latents) on the same training data as our main model (denoted unPIC-pixels in Table 4 or simply unPIC elsewhere). We use the VAE from Stable Diffusion 1.4 to encode all CROCS pointmaps, and source and target images. The VAE compresses images from 256x256x3 down to 32x32x4. We then train a diffusion model maintaining a similar architecture to unPIC-pixels. We concatenate latents for $K=8$ views to form a 4x2 grid as we did in the image case. Although the dimensionality of the latent grid is smaller, we keep all architectural hyperparameters the same, allowing unPIC-latents a slight advantage.
Test datasets. In addition to our 50k held-out examples of Objaverse XL, we evaluate on three out-of-distribution datasets: (i) Google Scanned Objects [8], (ii) Amazon Berkeley Objects [4], and (iii) the Digital Twin Catalog [36]. Some of the baselines may have trained on these test datasets—for instance, One-2-3-45 was trained on ObjaverseXL (we use the standard Zero123-XL checkpoint).
Metrics. To measure novel-view prediction accuracy, we compute a number of image-quality metrics such as PSNR, FID [15], LPIPS [56], and SSIM [48]. In addition, we focus on evaluating:
  • Geometric accuracy: we compute the IoU (Intersection over Union) of the 2D mask of a predicted novel view w.r.t. the expected mask. This metric is invariant to lighting or texture.
  • Multiview consistency: we compute CLIP embeddings (using the open-source ViT L/14 336px model) for all K generated images in a given example, then compute the mean of the (K x K) pairwise distance matrix of the embeddings. This metric (abbreviated CLIP-MV) avoids any comparison with ground-truth images, and merely probes the internal consistency of a set of images. All reported CLIP-MV numbers are 1e-4.
When computing metrics, we sample a random target from the set of 7 novel views for each example (except for CLIP-MV, where all novel views are used). This ensures the computed statistics are independent (particularly because FID relies on second-order statistics). We do not repeat any object (as in training) with an alternative source view. We run all models, including ours, on A100 GPUs, and downsize all images to 256x256 for comparison.
Results. Table 4 shows that unPIC beats all baselines on PSNR and SSIM. Notably it outperforms them substantially on the shape-focused IoU metric. We examine this advantage qualitatively in Figure 4. Although all methods tend to produce plausible looking images, the generated poses and shapes can be completely arbitrary (Figure 5). Being trained on Objaverse and ObjaverseXL alone, unPIC underperforms on the FID score in comparison to methods which likely benefit from ImageNet pretraining. Nevertheless, its LPIPS scores establish the perceptual quality of its outputs.
Figure 5: While all models produce plausible images, their shapes and poses can be off. We compare masks with the ground-truth.
We find a common failure mode that affects shape and pose prediction for CAT3D and OpenLRM: despite specifying a constant camera height, OpenLRM appears to level and raise the camera (w.r.t. the ground plane) as it moves around the object (see fig. 2 in the LRM paper [18]). The issue appears consistently in OpenLRM outputs, and intermittently in CAT3D’s as well. It likely stems from how the training data was generated for those models: the orientation of each object was likely reset before capturing images. As a result, CAT3D ends up modeling the stochasticity of the object’s initial orientation; OpenLRM, being a deterministic model, overfits to it. At inference time, these methods randomly guess an orientation, so they may assume the object is tilted w.r.t. the world’s up-direction. This leads to an apparent camera-control failure, where the camera appears to raise or lower itself as it moves around the object. It is especially undesirable for assets placed on a planar surface, as the whole surface appears to tilt while the camera orbits the object. In our case, we avoid this issue by ensuring the object’s vertical axis is always aligned with the world’s.
6 Discussion
Figure 6: Diversity of unPIC outputs from an ambiguous image (top-left). All samples show a 90-degree rotation from the source view, and are shown here without cherry-picking.
Figure 7: In-the-wild inference on natural images from two datasets: GQA [20] and Objectron [1]. As in Figure 1, the first column shows the source image; columns 2-5 show four of eight predicted CROCS pointmaps; and columns 6-9 show corresponding novel-view predictions. unPIC generalizes well beyond the digital assets it was trained on.
Multiview synthesis is central to image-to-3D modeling. A sparse but targeted set of novel views can serve as an ”anchor” for downstream generation of more fine-grained views, an arbitrary target pose, or an explicit 3D reconstruction. This underscores the importance of improving existing multiview priors, which tend to be image-focused and geometry-free.
There are several advantages of our hierarchical approach: first, it allows maximizing the size of the prior and decoder models, as they can be trained separately; second, we can visualize and interpret the intermediate features—for instance, visualizing CROCS could reveal whether 3D reconstruction errors arise due to geometry or texturing; and finally it prevents a collapse of output diversity in the prior—this would be much likelier if the modules were cotrained, as the prior would be spurred to output only the optimal intermediate representation for the final RGB loss. See Figure 6 for an example of diverse 3D shapes predicted from a single image.
We posit that unPIC’s main advantage is being geometry-driven. This was evidenced by the geometric accuracy and multiview consistency of unPIC’s novel-view predictions in Section 5. To provide further evidence that decoupling geometry and appearance supports generalization, we run unPIC on in-the-wild images far from its training distribution (see Figure 7). These results suggest that unPIC could be useful for a variety of downstream 3D applications, and help ensure that downstream tasks are well-posed.
Limitations & Future Work. A salient shortcoming is that we do not model scene backgrounds. This may be feasible using CROCS for room-like scenes where the background geometry is uniform, and can be scaled to a unit cube. We are currently also limited to training on images of 3D assets rendered at controllable poses, rather than in-the-wild images. This could be eased by using pretrained 3D models (e.g., NeRFs) of real scenes to render images and depth maps, which then could be used to estimate CROCS.
A straightforward extension of our study would be to allow multiple random source views rather than one. One of them could be designated the primary view. It would be interesting to compare the model’s performance in two cases: when it infers the correspondence between the source images, or when they are accompanied by camera pose annotations to avoid correspondence ambiguity.
While our ablation comparing the latent-diffusion implementation of unPIC with the pixel-diffusion-based implementation did not suggest any immediate advantage for unPIC-latents, this was likely due to the fact that the pretrained VAE did not offer a suitable latent space to encode CROCS pointmaps. It would be worth fine-tuning a VAE or training a geometry-focussed latent space.
The unPIC framework is not limited to one prior, e.g., it could be fruitful to predict multiview texture features to condition the decoder. It may also be possible to enhance multi-object handling by breaking CROCS into two components: a scene-level pointmap like ours (that colors objects based on their positions in the scene), and a pointmap that colors objects individually, capturing each object’s shape independent of its position in the scene.
Conclusion. We introduced the unPIC framework for multiview synthesis from a single image via multiview features. Our choice of intermediate features, CROCS, encodes scale-free geometry and point-to-point correspondence across views. We showed it is an effective choice over several geometric and non-geometric alternatives. It is predictable irrespective of the camera pose of the source image, and supports learning and generalization across examples.
Our two-tier system helps address the under-specification of image-to-3D tasks by allowing probabilistic outputs at each level. We encourage multiview consistency at each level using explicit geometric supervision and control. Our approach outperforms multiple competing alternatives on a variety of metrics, especially in terms of shape and pose accuracy. unPIC learns shape priors that generalize well beyond the digital assets it was trained on, paving the way for more geometry-driven novel view synthesis.
Appendix A Layout
The appendix is organized as follows: we describe CROCS in detail in Appendix B, discussing in particular the geometrical properties which led to our algorithmic choices. We then present additional qualitative comparisons and quantitative analyses in Appendix C. We break down unPIC’s hierarchical error in Section C.2. Finally, we describe all training and evaluation details for unPIC in Appendix D.
Appendix B Geometric Subtleties
B.1 CROCS Rescaling
Recall that we pre-render multiview NOCS images using a fixed reference frame. This reference frame can be labeled using the source-view camera position $(\theta,\phi,r)=(0,0,1)$ that captures the object in its default creator-intended pose at a default camera distance. At training time, we would like to set the source camera to arbitrary locations to ensure a rich training distribution. To change the reference frame from $\theta=0$ to a new source-camera location $\theta^{\prime}$ (we will treat the other parameters in Section B.2), we simply rotate the ground-plane axes of all pre-rendered NOCS maps using the SO2 rotation matrix $[[\cos\theta^{\prime},-\sin\theta^{\prime}],[\sin\theta^{\prime},\cos\theta^{\prime}]]$. The ground-plane axes are the Red and Green channels of the NOCS maps in our case.
Our on-the-fly NOCS-to-CROCS canonicalization comes with a challenge, arising from the fact that the object was scaled to fit the NOCS cube at $\theta=0$. When we rotate the RGB reference cube to follow the primary camera to $\theta^{\prime}$, then the object may no longer be bounded by the rotated cube. See Figure 8 where we visualize this issue for a fixed object scaled to fit the reference cube at $\theta=0$.
In the upper row of Figure 8(a), we get some CROCS values outside the range $[0,1]$, because the object exceeds the boundaries of the $[0,1]^{2}$ CROCS reference square. Some values of $\theta^{\prime}$ such as 45, 135, 215, and 305 degrees are affected the most, because they lead to the greatest object overhang. This is not ideal—the scale of CROCS values which the model needs to predict depends on the reference frame. Since the reference frame (i.e., source camera position) is not actually supplied to the model, the model can only learn to predict the variable scale by overfitting. We aim to avoid any dependence on the reference frame. To this end, we introduce a CROCS rescaling operation to eliminate the dependence (Figure 8(a), lower row).
Taking a concrete example, Figure 8(b) shows a cube-like object from different angles, along with the pre-rendered NOCS images (second row), and canonicalized CROCS maps without and with rescaling (third and fourth rows). Here are two observations from this figure: 1) The front-facing edge of the cube, visible at $\theta^{\prime}=45$ as yellow in the second row, disappears in the third row because the CROCS values exceed the $[0,1]$ range (and the color map is defined on the same range). If we rescale the values in the third row, we see the edge reappears in the fourth row. 2) The maximum overhang (in the third row) occurs at $\theta^{\prime}=45$, where the range of observed $y$ values expands to $[-0.2,1.2]$. The length of this expanded interval is nearly $\sqrt{2}\approx 1.41$, which can also be deduced geometrically—it equals the length of the diagonal of the fixed square in Figure 8(a).
Geometrically, we expect no object overhang at all if a given object is bounded by a sphere of diameter 1 (a tighter bound than the unit cube). We believe a lot of objects fall in this category. For this reason, our model does reasonably well even if we train it to predict CROCS without rescaling. But we observed a noticeable gain in the prior’s performance on CROCS with rescaling. The latter setting removes the reference-frame-dependent scale that the model needs to predict in some cases (objects like a cube).
Note that for a given object, the rescaling needs to be performed across all multiview pointmaps (K=8 target views in our case) simultaneously, since any given view only shows a particular facet of the object. (A single pointmap may entirely miss the variation along its depth component.) Our rescaling is reminiscent of the pairwise pointmap renormalization performed by DUSt3R (see Eq. 3 in [47]). All the results in this work use CROCS with rescaling.
(a) The general case (simplified to 2D). Top row: An object that fits the NOCS reference square exactly at $\theta=0$ is no longer bounded by the CROCS reference square at $\theta^{\prime}\in\{15,30,45\}$. Due to object overhang, some CROCS values lie outside $[0,1]$. The extent of the overhang depends on $\theta^{\prime}$. Bottom row: We rescale CROCS based on the observed range of multiview values for any $\theta^{\prime}$. This ensures the object is once again bounded tightly by $[0,1]^{2}$.
(b) A concrete example in 3D. We show the minimum and maximum NOCS/CROCS values observed above each pointmap image. At $\theta^{\prime}=45$, we see that the front-facing edge of the cube appears squashed without rescaling, because its X and Y values lie outside the range [0, 1].
Figure 8: CROCS rescaling. We examine the effect of rotating the RGB reference cube used to paint the object’s surface, following the source camera as it moves around a fixed object. $\theta$ (the azimuthal angle) denotes the default camera position, while $\theta^{\prime}$ denotes a new position. We consider the largest possible objects—a square in 2D or cube in 3D.
Figure 9: CROCS when varying the camera elevation angle $\phi$. As in Fig. 3, the large camera denotes the source view; the smaller cameras denote 3 of 7 novel views. The white arrow denotes the normal direction w.r.t. the camera plane (drawn as a white circle connecting the cameras). The colored arrow denotes the normal of the CROCS color cube. Left: at the default angle $\phi=0$, each camera is pointed at a particular face of the RGB color cube, and has a consistent color distribution across examples. Middle: when the source camera is raised to $\phi^{\prime}$, we could hypothetically rotate the color cube by the same degree to ensure the source camera still points at the same face of the cube. The issue is—tilting the color cube toward the source camera tilts it away from the opposite camera, and also rotates the colors seen by the remaining target cameras (by $\phi^{\prime}$ and $-\phi^{\prime}$ in their respective image planes). The color shifts experienced by the four cameras are clearly inconsistent. Right: If we choose not to tilt the cube through $\phi^{\prime}$, all cameras undergo a small and consistent shift in the distribution of colors toward the up-axis color (blue). If the model can infer $\phi^{\prime}$, it can also predict the shift. We follow the Right approach.
B.2 No Canonicalization for Camera Elevation or Distance
In Section B.1 we discussed how to canonicalize NOCS when switching the source camera from $\theta$ to $\theta^{\prime}$. Here, we discuss the remaining camera extrinsic parameters. Since all cameras are oriented to face the center of the object, we do not need to consider their rotations. In fact, we do not need to canonicalize for changes in the camera distance $r$. It only serves as a zoom parameter to augment the training data.
The only remaining parameter is the elevation angle $\phi$, which determines the source-camera height. Recall that the target cameras are also placed at the same camera height. In fact, $\phi$ determines the difficulty of the multiview prediction problem. For instance, at $\phi=90$ (a straight-down view of the object), the source and target camera locations converge to the same point—the apex of the hemisphere. In this case, the prediction task is reduced to rotating the object in the image plane.
Say we raise/lower the source camera from $\phi$ to $\phi^{\prime}$ (Figure 9). This would reveal more/less of the object’s upper/lower surface, and hence more/less of the NOCS color representing the up-axis (Blue in our case). We could potentially correct for this by tilting the reference cube so the source camera is pointed at the same cube face as before at $\phi$. This would help maintain the same color bias for the source camera. However, the target cameras would no longer point at the same cube faces respectively—rather, their distribution of colors would shift (toward the up-axis/away from it in the case of the opposite target camera) or rotate (in the image plane for other target cameras). Consequently, the distribution shifts in the CROCS colors would be inconsistent across target cameras.
Based on this geometrical observation, we choose not to canonicalize for changes in the camera elevation $\phi^{\prime}$. This has the effect of leaving it to the model to infer the camera elevation from the source image (implicitly in contrast to One-2-3-45 [30]). It differs from our treatment of $\theta^{\prime}$, which leaves the model oblivious to the azimuthal rotation of the cameras.
Our decision not to canonicalize for $\phi$ is forced by the design choice of our target camera poses, which follow the camera height of the source camera. An alternative choice would be to tilt the camera plane and the CROCS cube together, leaving the cameras at different heights in Figure 9-Middle [^1] This alternative is equivalent to using a stationary multiview camera rig, but orienting the object randomly before taking pictures, as in CAT3D [12], likely LRM [18], and datasets like YCB [3]. Changing the object’s orientation decouples the object’s up-axis from the world’s up-axis. It can leave the object in unnatural, physically unstable poses.. Our choice of target poses is based on how humans perceive and reason about objects. We tend to see side or top-down views of level objects. On mental rotation tasks (e.g., when asked what an object looks like from “behind”), we tend to imagine a rotation of the object either in the image plane, or in depth about the world’s vertical axis [42, 44]. We find it harder to imagine a potentially bottom-facing view. In addition to aligning with human perception, modeling object rotations at typical views/poses (e.g., top-down or side angles) could also be more useful for downstream applications.
Appendix C Additional Results
C.1 Comparisons
Figure 10: Additional qualitative comparison on Objaverse-XL holdouts.
Figure 11: Additional qualitative comparison on Objaverse-XL holdouts.
Figure 12: Additional qualitative comparison on Objaverse-XL holdouts.
Figure 13: Additional qualitative comparison on Objaverse-XL holdouts.
We present further qualitative results comparing unPIC with CAT3D and One-2-3-45 (the top-two competing methods) in Figures 10, 11, 12 and 13.
Besides its multiview inconsistency, One-2-3-45 also has the flaw that it rescales any given object to a canonical size in the image plane. This is visible from the model’s source-pose output (top-left cell in each 2x4 output grid); it is resized compared to the source-pose outputs of other methods. Our approach can cope with arbitrarily sized objects in images despite the fact that it relies on a scale-free 3D shape representation (CROCS).
Note that both CAT3D and our method predict gray values at background pixels. But unPIC produces RGBA outputs rather than RGB—the alpha mask hides the gray background pixels after compositing. We compute masks using posthoc background removal for CAT3D. The remaining methods (One-2-3-45 and OpenLRM) output RGB images with white background pixels.
C.2 Decomposing unPIC’s hierarchical error
In Sec 4.3 we reported metrics from the full hierarchical setup, i.e., on the final output of the prior and decoder stacked together. Since we train the prior and decoder modules separately, we can further analyze the following components of the total error:
  1. 1.
    The error of the prior alone. This describes the geometrical inaccuracy (when the intermediate representation is CROCS) in predicting the 3D shape and pose of a given object.
  2. 2.
    The error of the decoder when it is fed ground-truth CROCS rather than the output of the prior. This describes the difficulty of rendering the object (e.g., predicting the object’s texture) at novel views when extrapolating from a single source image.
We report these components along with total hierarchical error in Table 5. We find that the prior’s error (column 1) is nearly twice that of the oracle decoder (column 2). This highlights the importance of getting the geometry correct. Improving the geometrical prior would help reduce the total error (column 3) more than improving the final rendering (i.e., the decoder module). With geometry predictions as good as the ground-truth CROCS, our full multiview-synthesis error could be 4-5x lower (depending on the dataset).
Table 5: A breakdown of unPIC’s hierarchical error (256x256). We report Mean Square Errors (1e-3) for the prior module, the decoder module when using ground-truth CROCS, and the full error running stacked inference. The MSE is calculated on all 7 novel views.
Image
$\rightarrow$ CROCS
GT CROCS
$\rightarrow$ Pixels
Image
$\rightarrow$ CROCS
$\rightarrow$ Pixels
ObjXL 4.58 2.84 10.52
ABO 3.46 1.65 7.60
GSO 2.69 3.09 7.63
Appendix D More Training and Evaluation Details
Dataset. We use the following sets of assets from Objaverse: 1) the LVIS subset expanded to 83k examples from the same categories. We use object type labels predicted with high confidence from [24] to expand the LVIS subset. 2) The KIUI subset  [^2] https://github.com/ashawkey/objaverse_filter comprising 101k additional assets. 3) From the Objaverse-XL alignment subset [6], we used a combination of GLB, OBJ, and FBX assets after filtering out certain (a) terrains and HDRI environment maps, (b) layouts and rooms, and (c) textureless FBX assets that rendered as pink. In total, we used 620k assets (Objaverse and Objaverse-XL) for training. In addition, we held out 50k assets randomly sampled from Objaverse-XL after the filtering step.
To render NOCS images, we adapted a script from BlenderProc [^3] https://github.com/DLR-RM/BlenderProc/ [7] which creates a special NOCS material for the surface of a given object. We also use their Blender settings (the CYCLES engine with 1 diffuse bounce, 0 glossy bounces, and 0 ambient occlusion bounces) for rendering NOCS. We export them in the EXR format to ensure linear-light values, preserving the continuity of NOCS values on the object’s surface.
Model. We assume an RGB source image, but predict RGBA CROCS and RGBA output images. We train both models with a conditioning dropout probability of 0.05. This opens the possibility of using classifier-free guidance (CFG). For all evaluations in the paper, we use a fixed CFG value of 2.0. The prior model is normally conditioned on the timestep and the source image. We apply CFG to the source image. The decoder model is additionally conditioned on the prior’s output CROCS. We only apply CFG to the source image, as the decoder tracks CROCS pointmaps closely even without CFG.
The architecture for the prior and decoder has just under 150M parameters. It takes about 6 minutes to run the prior and decoder each for 2000 denoising steps, with classifier-free guidance, on an A100. As efficiency wasn’t a core goal for this work, we expect more efficient sampling strategies and better architectures could certainly help reduce the runtime. We use an exponential moving average of the model parameters for evaluation, although this is not essential.