### v3

# How to Spin an Object: First, Get the Shape Right

Abstract
Image-to-3D models increasingly rely on hierarchical generation to disentangle geometry and texture. However, the design choices underlying these two-stage models––particularly the optimal choice of intermediate geometric representations––remain largely understudied. To investigate this, we introduce unPIC (undo-a-Picture), a modular framework for empirical analysis of image-to-3D pipelines. By factorizing the generation process into a multiview-geometry prior followed by an appearance decoder, unPIC enables a rigorous comparison of intermediate geometry representations. Through this framework, we identify that a specific representation, Camera-Relative Object Coordinates (CROCS), significantly outperforms alternatives such as depth maps, pretrained visual features, and other pointmap-based representations. We demonstrate that CROCS is not only easier for the first-stage geometry prior to predict, but also serves as an effective conditioning signal for ensuring 360-degree consistency during appearance decoding. Another advantage is that CROCS enables fully feedforward, direct 3D point cloud generation without requiring a separate post-hoc reconstruction step. Our unPIC formulation utilizing CROCS achieves superior novel-view quality, geometric accuracy, and multiview consistency; it outperforms leading baselines, including InstantMesh, Direct3D, CAT3D, Free3D, and EscherNet, on datasets of real-world 3D captures like Google Scanned Objects and the Digital Twin Catalog.
1 Introduction
Recovering 3D appearance from a single image is a hard, underspecified problem [11, 25]. Models that provide 3D coverage by generating novel viewpoints of a given scene must meet two goals: one, maintain 3D consistency across the generated novel views; and two, realize diverse 3D instantiations of the given 2D input. Multiview diffusion models [21, 10, 59, 53], which sample multiple novel views together, have achieved considerable success at this task.
[…]
Despite its well-motivated foundations, hierarchical 3D generation (image $\rightarrow$ 3D geometry $\rightarrow$ 3D appearance) remains understudied. Commercial approaches like Hunyuan3D [58] and CLAY [56] use specialized 3D geometry encoders to represent object shapes, with little analysis of design choices. Furthermore, there is limited insight into (a) the value of a hierarchical pipeline, and (b) the optimal choice of intermediate representations. To study these questions, we introduce a modular image-to-3D framework called unPIC (undo-a-Picture). We make two assumptions to facilitate systematic study: we use dense, image-like representations of 3D shape (e.g., depth, segmentation, or pointmaps), and off-the-shelf VAEs (from Stable Diffusion) to encode geometry and appearance.
We find a particular choice of intermediate representation outperforms others: Camera-Relative Object Coordinates. CROCS encodes the 3D coordinates of all points in a scene within a unit cube oriented toward the source camera. We find in Section 4.1 that (i) CROCS is an effective conditioning signal for the second-stage appearance decoder compared to alternatives such as depth, other pointmaps (NOCS), or pretrained visual representations (DINO or CLIP); and (ii) CROCS is also easier to predict for the first-stage geometry prior. Putting the two stages together, we show that unPIC using CROCS achieves (iii) superior novel-view accuracy and multiview consistency than strong multiview diffusion baselines including CAT3D, EscherNet, One-2-3-45, and Free3D (Section 4.2).
Conveniently, CROCS enables feedforward, direct 3D generation of a point cloud without a separate reconstruction stage. We show that (iv) direct 3D reconstruction via CROCS is more accurate than SoTA baselines Direct3D or InstantMesh (Section 4.3). Finally, we (v) ablate unPIC’s hierarchical formulation in Section 4.4, showing a significant benefit to predicting the shape of an object before filling in appearance-level details.
Figure 1: The unPIC architecture. A prior predicts multiview geometric features from a single source image. Our choice of features, CROCS, establishes point-to-point correspondence across views. The predicted CROCS images serve as a blueprint for the decoder module, which generates the fully textured target-view images, spinning the object around its vertical axis. We show 4 of 8 views here for clarity. All images are encoded as latents for diffusion. Once decoded, the output images can be directly converted to the vertices and vertex colors (respectively) of a 3D point cloud.
2 Related Work
2.1 Background
Classic novel view synthesis (NVS): Scene‑specific methods such as NeRF [28] and 3D Gaussian Splatting [20] fit explicit geometry but require multiple consistent input views. When only one image is available, these methods need a learned prior to predict what the scene looks like from alternative views. Such priors include pixelNeRF [54] and LRM [15]. They rely on spatially arranged representations to enable feed-forward NVS, but are deterministic, and suffer from averaged/blurry reconstructions.
Diffusion-based 3D generation: Early text-to-3D methods such as DreamFusion [32] and Score Jacobian Chaining [41] introduced the use of diffusion priors for iterative 3D generation. But they optimize per scene and can suffer from multi-face (Janus) artifacts. Subsequent image-to-3D models like Zero-1-to-3 [27] and ZeroNVS [35] train 3D-aware diffusion models from single images, enabling feed-forward inference, typically followed by an optimization step to produce a final 3D representation. This “generate then optimize” paradigm has become a dominant strategy for 3D reconstruction.
Multiview diffusion for NVS: Recent methods such as EscherNet [21], Free3D [59], ReconFusion [48], and CAT3D [10] denoise K views simultaneously to exchange information across images and promote consistency. Nevertheless, their lack of geometric grounding can result in generated images that are individually plausible but collectively represent an impossible 3D object. Efforts to mitigate this, such as 3DiM’s stochastic conditioning [46], still lack direct geometric control.
2.2 Motivation
Geometry as an input: ControlNet [55] popularized the use of pixel-aligned hints to condition and guide a diffusion model. Other successful applications include motion-brush animation [29] and object-pose-conditioned scene/layout generation [16, 50]. These non-hierarchical approaches are especially tempting given that diffusion models can be trained with conditioning dropout to enable classifier-free guidance [14]. At inference time, the model can cope with not having any geometric information, and still sample from the unconditional distribution.
Geometry as an output: Pretrained diffusion models yield useful 3D information even when they are entirely appearance-focused [8, 7]. More geometry-focused models like He et al. [12], Ke et al. [19], Fu et al. [9] output dense annotations such as depth and surface normals, but typically annotate monocular images rather than predict novel views. They lack geometric consistency when applied to different views of a scene. We show that our pointmap-based representation (CROCS) facilitates explicit, 3D-consistent geometry prediction.
Hierarchical generation: unPIC decouples geometry and appearance by first predicting novel-view geometry via a dedicated prior, then conditioning an appearance decoder on the result. This is similar to Motion-I2V’s explicit motion generation to condition a motion-guided video model [37]. A hierarchical image-to-3D pipeline produces realistic, diverse 3D shapes without manual effort. A separation of concerns between the geometry prior and appearance decoder ensures that each stage remains specialized, promoting higher predictive diversity than end-to-end models.
2.3 Current Approaches
Geometry‑supervised image-to-3D. Methods such as One‑2‑3‑45 [26] (SDF via cost volumes) and InstantMesh [53] (triplane to mesh with depth/normal supervision) incorporate explicit geometric signals and return explicit surfaces. Closer to our work, Bolt3D [38] and CRM [45] use pointmaps as their geometric representation; however, geometry and appearance are predicted jointly (Bolt3D) or appearance-first (CRM), rather than unPIC’s geometry‑first approach. Two recent approaches, TRELLIS [51] and CLAY [56], also use a geometry-first approach, but rely on specialized 3D VAEs rather than our 2D-pretrained VAE (from Stable Diffusion 1.4). A separate line of work focuses on untextured mesh generation (e.g., CraftsMan3D [24], Direct3D [49]), emphasizing shape quality.
Pointmaps: DUSt3R [43] predicts dense pointmaps to establish 3D correspondence between images, but does not model unseen views. VGGT [42] proposes viewpoint‑invariant pointmaps with per‑scene normalization. SpaRP [52] uses a diffusion model to estimate NOCS pointmaps [40] for the given input views; in the single-image-to-3D setting, SpaRP would predict a single NOCS image to annotate the source view only, as the model does not explicitly predict novel-view geometry.
SweetDreamer [23] is closer to unPIC in its use of ‘Canonical Coordinate Maps’ to predict novel-view geometry, but still relies on the NOCS-assumption that “all objects within the same category adhere to a canonical orientation”; it further relies on optimization (score distillation) to produce the final 3D reconstruction. Our CROCS differs by being camera‑relative unlike SweetDreamer, and normalized consistently to $[0,1]$ unlike VGGT. CROCS yields predictable statistics across target viewpoints, and enables direct point‑cloud extraction during view synthesis.
3 Method
3.1 Hierarchical generation
unPIC is a hierarchical generative model that takes a 2D image and predicts a set of target novel-view images. Given a single view of an object, unPIC comprises: (i) a geometry prior that infers representations of the object’s 3D geometry, and (ii) an appearance decoder that transforms all multiview geometry predictions into textured images of the subject. The two modules are trained independently. At inference time, we run the modules in a stacked fashion: the decoder takes the prior’s outputs and the source image as inputs.
This specialization and separation of concerns has two advantages: one, it helps maximize the diversity of our samples at each stage. If the prior and decoder were trained end to end, the prior would need to output the most optimal representations for the decoder, and those could be texture-specific rather than geometric. A second advantage arises from the fact that the appearance decoder is only ever trained with ground-truth geometry: it adheres closely to the geometric conditioning.
3.2 Multiview Diffusion
Both the prior and decoder modules are implemented with identical multiview diffusion architectures, building on recent work [48, 21, 59, 10] that showed the benefit of predicting all K target views simultaneously. We found that tiling the views to form a single superimage works well; we tile them clockwise to ensure each view is adjacent to its closest neighbors in camera space (see Fig 1). Our U-Net architecture follows [10], ensuring information exchange at two levels: between adjacent views using standard convolutions, and globally using self-attention across the superimage. Our implementation models K=8 views in a 2x4 grid. K=8 is sufficient to capture object geometry and appearance, and allows enough viewpoint overlap to teach the models to maintain multiview consistency, as further motivated in Appendix B.3. Since the target views are defined w.r.t. the source camera pose and remain fixed to produce a $360^{\circ}$ spin, our model does not require explicit camera poses as input. The prior and decoder modules are trained from scratch, exclusively on Objaverse and ObjaverseXL assets.
We follow the Latent Diffusion approach [34], using a VAE to encode the K views independently before tiling them to form a superimage. We fine-tune the Stable Diffusion 1.4 VAE separately on CROCS and RGB images from our training data under identical training conditions. While both VAEs start with similar PSNR scores ($\sim$37-38), the CROCS VAE’s score goes up to 51 after fine-tuning; the RGB VAE’s score improves slightly to 38. The CROCS VAE’s boost in performance comes at a lower KL representational cost than the RGB VAE’s (see Table 7 for detailed metrics). This is intriguing given that both VAEs started from the same pretrained checkpoint, which was never originally trained on CROCS images. We conclude that it is easier to reproduce CROCS pointmaps, which vary smoothly from pixel to pixel, than textured RGB images. This is yet another motivation for unPIC’s approach of predicting geometry before appearance.
Our full framework can be described as follows: Let $z^{g}_{k}$ denote the geometry latent for the $k^{th}$ view (encoded from a CROCS pointmap) and $z^{a}_{k}$ denote the appearance latent (encoded from an RGB image). We assume both kinds of latents are standardized by their empirical mean & standard deviation. They are then tiled to form the target superimages for geometry ($Z^{g}$) & appearance ($Z^{a}$):
$Z^{g}=\text{tile-clockwise}(\{z^{g}_{k}\}_{k=1}^{K})$
$Z^{a}=\text{tile-clockwise}(\{z^{a}_{k}\}_{k=1}^{K})$
Both the prior and decoder are conditioned on an input superimage, $X^{a}$, which is constructed from a masked set of K appearance latents. During training, we sample the index of a single source view $k^{*}\sim Uniform({1,…,K})$. The input latent for this view, $x_{k^{*}}$, is set to its ground-truth appearance latent, $z_{k^{*}}$, whereas the remaining K1K-1 latents are replaced by 0. While we train on only one input view, randomizing its position teaches the model to process arbitrary input configurations at inference time. The resulting input latents, $\{x^{a}_{k}\}_{k=1}^{K}$, are tiled clockwise to form the final superimage:
$X^{a}=\text{tile-clockwise}(\{x^{a}_{k}\}_{k=1}^{K})$
Finally, the unPIC inference pipeline can be written as:
$\hat{Z}^{g}\sim f_{\text{prior}}(\cdot|X^{a})\quad\text{and}\quad\hat{Z}^{a}\sim f_{\text{decoder}}(\cdot|X^{a},\hat{Z}^{g})\quad$
3.3 CROCS: Camera-Relative Object Coordinates
Geometry underpins unPIC’s hierarchical approach: we predict the 3D shape of an object before decoding its multiview appearance. To do so, we propose a dense pointmap representation called Camera-Relative Object Coordinates (CROCS). CROCS is a scale-free representation of geometry: the object/scene is first uniformly scaled to fit a 3D unit cube, so all spatial coordinates are in [0, 1]. These can therefore be interpreted as RGB colors and rendered as images, which maps naturally onto our multiview framework. While a single CROCS image serves as a dense annotation for a target novel view, a set of CROCS images taken together form the vertices of a 3D point cloud.
CROCS can be seen as a variation of Normalized Object Coordinates [40]. Instead of orienting the color space w.r.t. the object’s class-canonical pose, CROCS orients the scene geometry w.r.t. a given source camera. CROCS is also related to VGGT’s view-invariant pointmaps [42], but those are scaled by a per-scene normalization factor, and not actually bounded.
NOCS vs CROCS. While NOCS relies on a class-canonical orientation to establish correspondence between object instances, CROCS provides a pose-free, scene-level representation (see Figure 2). By orienting the coordinate system relative to the source camera rather than the object’s class, CROCS eliminates the need for object segmentation or identification. This camera-relative alignment ensures that color distributions remain predictable across target viewpoints—for example, the CROCS image for the source-camera view, corresponding to the front face of the CROCS cube, is colored white in the top-right-near corner and black in the bottom-left-far corner. This statistical predictability makes CROCS ideally suited for predicting novel views at regular intervals, as the model learns to associate specific camera aspects with biased color patterns.
We provide a precise mathematical description of CROCS in Appendix A.1. We also address two geometric subtleties in working with CROCS in Appendix A.2—the first one is about renormalizing pointmaps each time the source camera is moved to ensure they are still bounded in $[0,1]$. The second one explains why the CROCS cube remains axis-aligned, sitting squarely on the ground-plane, when the source camera is moved higher or lower (changing its elevation $\phi$).
Figure 2: Choice of pointmaps. Left: source images derived by moving the camera. NOCS: objects are normalized and colored individually based on a class-canonical pose. The same semantic part receives the same color across instances, e.g., a right arm is always blue. CROCS: objects in the scene are normalized and colored collectively. Each pixel encodes normalized $(x,y,z)$ coordinates. The source camera sets the orientation of the coordinate system (shown as the RGB cube, fixed across examples). Any target viewpoint around the cube has a predictable color pattern.
4 Experiments
We provide empirical results on our choice of representation in Sec 4.1; the quality of our novel-view synthesis in Sec 4.2; the accuracy of our 3D geometry in Sec 4.3, and finally an ablation of our hierarchical design in Sec 4.4.
4.1 Choice of Intermediate Representations
To motivate geometric representations and the use of CROCS specifically, we run two preliminary experiments: one to assess what features are predictive of novel-view images, and another experiment to assess what features are predictable from a single source image.
Predictiveness. We train a primitive version of unPIC’s decoder module to predict a single novel view of an object at a fixed target rotation ($\theta=90$ degrees). The model is conditioned on a source image, as well as a ground-truth annotation of the target novel view. The type of annotation is varied across training runs to assess the usefulness of different representations. The annotations range from geometric (like depth maps and CROCS), image-space maps (alpha masks), through feature maps (like DINOv2 [30] and CLIP [33]). All annotations are treated as images—they are upscaled (if required) and concatenated to the input image before it is denoised. We run these experiments at 64x64 image size (to feasibly upscale the high-dimensional CLIP and DINOv2 features), and using pixel-based diffusion to avoid ambiguity about how to encode the annotations. Table 1(a) shows that CROCS is the most helpful for this task.
Table 1: We analyze the effectiveness of CROCS for novel view synthesis (top) and its predictability from a single image (bottom) versus other geometric representations.
(a) Does CROCS help predict novel views? We predict a $90^{\circ}$ object rotation using varied target image annotations. We report pixel MSE and stds. over 3 training runs.
Annotation Type MSE (1e-3)
Source image only 19.042 $\pm$ 0.722
+ CLIP ViT L/16 6.059 $\pm$ 0.448
+ DINOv2 B/14 5.734 $\pm$ 0.815
+ Alpha mask 8.169 $\pm$ 0.392
+ Depth map 5.580 $\pm$ 0.849
+ NOCS 5.318 $\pm$ 0.070
+ CROCS 4.914 $\pm$ 0.224
(b) Which multiview features are predictable from a single image? We report the pointmap Mean Square Error (1e-3) at K=4 and K=8 target views.
K Annotation Type MSE (1e-3)
4 NOCS 11.94
CROCS 1.21
8 NOCS 15.82
CROCS 3.92
Predictability. Having shown that CROCS improves the decoder, we now ask: can CROCS itself be predicted reliably? We train primitive versions of the unPIC prior to predict either NOCS or CROCS from a single source image. We experiment with $K=4$ or $K=8$ target views. We compare two choices of representations: (i) NOCS images in a static reference frame, determined by the default creator-intended pose for the given asset; and (ii) CROCS, where we change the reference frame based on the source camera. We find a 4x increase in predictability from using CROCS (see Table 1(b)). With CROCS, each target pose has a biased distribution of colors, consistent across objects. This makes the geometric feature space unambiguous, freeing up the model’s probabilistic capacity for the ambiguous task of predicting the unseen object/scene geometry.
4.2 Task 1: Novel View Synthesis
We compare unPIC against a diverse set of seminal and state-of-the-art baselines: CAT3D, EscherNet, Free3D, OpenLRM, InstantMesh (Large), and One-2-3-45 XL. The key conceptual differences are summarized in Table 2 and detailed further in Appendix B.7.
Test data. We sample 512 test objects from 4 different datasets, including our held-out subset of Objaverse XL, and three out-of-distribution datasets: (i) Google Scanned Objects [6], (ii) Amazon Berkeley Objects [3], and (iii) the Digital Twin Catalog [31]. The baselines may have trained on these test datasets—for instance, One-2-3-45 was trained on ObjaverseXL. For each test object, we sampled a random source and target view, and evaluated all models on the same pair.
Metrics. To measure novel-view prediction accuracy, we compute a number of image-quality metrics such as PSNR, FID [13], LPIPS [57], and SSIM [44]. In addition, we focus on evaluating:
  • 2D Geometric Accuracy: We compute the IoU (Intersection over Union) of the binarized mask of the predicted novel view w.r.t. the expected mask. This metric is invariant to lighting or texture.
  • Multiview Consistency: Though we attempted to run MET3R [1] to measure multiview consistency, we found it inadequate for our use case, even for ground-truth images (see Appendix B.6). Instead, we compute CLIP (ViT L/14 336px) embeddings for K=8 generated images per object, including the source image. Then, we compute the mean of the (K x K) pairwise distances between the embeddings. This metric (abbreviated CLIP MV) helps probes the consistency of a set of generated images without comparing to ground-truth images.
Results. unPIC consistently outperforms the baselines on all novel-view metrics (Table 3). The IoU metric provides an early sign of its geometric accuracy. On the CLIP MV metric, it is noteworthy that the only method which outperforms unPIC consistently, InstantMesh, is also geometry-supervised.
Table 2: Summary of different approaches for single-image-to-3D generation. P: Probabilistic/Generative; D: Deterministic; $F$ Recon: Feed-Forward Reconstruction; V-Gen $\rightarrow$ Opt: View Generation followed by per-scene Optimization; SVD: Single-View Diffusion (run repeatedly per target viewpoint); MVD: Multi-View Diffusion (generates views jointly).
Model Output Model Type Generative Core Intermediate 3D Rep. Direct Geometric Supervision
LRM D FF Recon - Triplane $\rightarrow$ NeRF No (Image loss only)
Free3D P V-Gen $\rightarrow$ Opt MVD Implicit (No 3D Rep) No (Image loss only)
CAT3D P V-Gen $\rightarrow$ Opt MVD Implicit (3D Attention) No (Image/Perceptual loss)
EscherNet P V-Gen $\rightarrow$ Opt MVD Implicit (MV Attention) No (Image loss only)
One-2-3-45 P V-Gen $\rightarrow$ FF Recon SVD SDF (from Cost Vol.) Depth
InstantMesh P V-Gen $\rightarrow$ FF Recon MVD Triplane $\rightarrow$ Mesh (FlexiCubes) Depth & Normals
unPIC (ours) P Direct 3D + V-Gen MVD (Hierarchical) Pointmaps (CROCS) Pointmaps (CROCS)
Table 3: Novel-View Synthesis: empirical comparison on 4 datasets.
Dataset Method PSNR $\uparrow$ IoU $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ CLIP MV (1e-4) $\downarrow$
Objaverse XL CAT3D 12 0.61 82 0.27 0.64 2.52
EscherNet 13 0.63 87 0.24 0.68 3.13
InstantMesh 13 0.62 76 0.25 0.67 1.89
Free3D 15 0.66 82 0.22 0.71 3.10
One-2-3-45 12 0.57 72 0.26 0.65 2.55
OpenLRM 13 0.60 94 0.24 0.71 2.38
unPIC (ours) 24 0.72 63 0.15 0.80 2.01
Google Scanned Objects CAT3D 19 0.64 87 0.26 0.72 2.98
EscherNet 14 0.72 76 0.26 0.66 2.78
InstantMesh 14 0.71 61 0.25 0.66 2.12
Free3D 14 0.70 93 0.25 0.68 3.63
One-2-3-45 13 0.64 60 0.26 0.65 2.74
OpenLRM 14 0.66 109 0.26 0.69 2.93
unPIC (ours) 23 0.77 62 0.17 0.77 2.33
Amazon Berkeley Objects CAT3D 18 0.59 82 0.29 0.67 2.34
EscherNet 11 0.65 73 0.30 0.59 2.35
InstantMesh 12 0.67 60 0.27 0.62 1.54
Free3D 12 0.69 73 0.27 0.63 2.92
One-2-3-45 11 0.61 50 0.28 0.61 2.22
OpenLRM 11 0.57 122 0.30 0.65 2.55
unPIC (ours) 25 0.88 40 0.15 0.78 1.74
Digital Twin Catalog CAT3D 18 0.70 83 0.23 0.71 2.68
EscherNet 14 0.74 75 0.23 0.68 2.37
Free3D 14 0.76 87 0.22 0.70 2.90
InstantMesh 14 0.74 58 0.22 0.68 1.51
One-2-3-45 14 0.74 53 0.21 0.69 1.98
OpenLRM 14 0.70 98 0.22 0.72 2.40
unPIC (ours) 25 0.90 46 0.13 0.82 1.55
Table 4: Ablation: we train a non-hierarchical version of unPIC without any geometry.
Dataset Method PSNR $\uparrow$ IoU $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ CLIP MV (1e-4) $\downarrow$
Objaverse XL unPIC ablation 20 0.63 89 0.23 0.75 2.07
Google Scanned Objects unPIC ablation 21 0.69 97 0.24 0.75 2.43
Amazon Berkeley Objects unPIC ablation 19 0.61 84 0.28 0.70 1.84
Digital Twin Catalog unPIC ablation 22 0.82 69 0.19 0.78 1.67
Table 5: 3D Geometric Accuracy: we compute the bidirectional Chamfer distance (1e-3) $\downarrow$ to compare point cloud completeness and accuracy. All methods take a single image as input. GSO: Google Scanned Objects. DTC: Digital Twin Catalog.
Method GSO DTC
EscherNet [21] 31.4 -
Direct3D [49] 27.1 -
Zero123++ & MeshLRM [47] 15.7 -
Direct3D (re-run) 8.94 5.99
InstantMesh (re-run) 6.83 6.62
unPIC (ours) 4.59 5.47
Figure 3: Novel views, baselines vs unPIC. The left column shows the source view, as output by each model. The models show different failure modes: (i) mutations of the original object (e.g., InstantMesh, EscherNet; annotated in blue); (ii) multi-view inconsistencies (e.g., One-2-3-45, Free3D; annotated in red); (iii) issues inferring the depth of the boat from the source image (e.g., CAT3D, OpenLRM); and (iv) pose inaccuracies in the novel views (e.g., InstantMesh gets the pose of the sail right, but the pose of the boat wrong; annotated in violet).
Figure 4: Points clouds, baselines vs unPIC.
4.3 Task 2: 3D Reconstruction
We derive point clouds from unPIC by simply combining the multiview outputs. The CROCS images provide the vertices, while the RGB images provide the vertex colors; we only need to mask out the background pixels from the output images. We run two competing methods, Direct3D and InstantMesh (Large), on our test images for comparison. To ensure the Chamfer distances are comparable, we subsample 8192 points from every point cloud. We also provide the numbers reported by EscherNet, MeshLRM, and Direct3D; our Direct3D re-run produced a similar score as the original paper. Table 5 shows that unPIC outperforms all methods. See Figure 4 for a qualitative comparison.
4.4 Ablating Hierarchical Generation
We train a version of unPIC without any geometry inputs or supervision, under identical training conditions. It is non-hierarchical as it takes a single RGB image and outputs the target novel views. In Table 4, we find that the ablated model performs worse than the full version of unPIC; nevertheless, it is comparable to CAT3D, which is another geometry-free multiview synthesis method.
5 Discussion
Figure 5: Diversity of outputs. Top row: various test images from DTC and ObjXL. Remaining rows: unPIC’s output at a fixed $90^{\circ}$ rotation across three RNG seeds, showing the range of outputs for the given conditioning.
Generative diversity. We qualitatively demonstrate the advantage of a hierarchical approach by sampling multiple outputs for a fixed input image. Figure 5 shows that unPIC’s novel views remain plausibly constrained while exhibiting appropriate variety across random seeds. A hierarchical pipeline like unPIC helps capture multimodal distributions in both geometry (varying structural interpretations of ambiguous inputs, e.g., the pomegranate or toy blocks) and appearance (texture variations, e.g., the white device).
Generalization and real-world utility. Having never seen a real-world pixel during its training, unPIC nevertheless generalizes to in-the-wild images (see Figure 23). This is clear evidence that it is possible to learn real-world shape priors from synthetic assets alone. While our focus was on using these priors for image-to-3D generation, unPIC could be repurposed for vision tasks, e.g., monocular or stereo depth estimation.
Limitations and failure cases. (i) A salient shortcoming is that we do not model image backgrounds. This may be feasible while still using CROCS where the background geometry is uniform, and can be scaled to a unit cube (e.g., room-like scenes). (ii) unPIC rotates some objects in the image plane rather than spinning them around the vertical axis, likely because it misinterprets the source image as an overhead view of the object. Since CROCS does not account for the source camera elevation, we could potentially condition unPIC on the elevation to prevent ambiguity. We provide an example of this failure in Figure 24. (iii) Natural images containing faces or multiple people can be out-of-distribution, as unPIC was trained on synthetic assets lacking such cases. We hope that expanding our training data, e.g. by combining multiple assets in the same scene, could offer one avenue to ease this limitation. (iv) At present, unPIC is also limited by the expressiveness of the SD-VAE: for example, it cannot represent any text. A richer representation space could help get over this limitation.
Conclusion. We introduced the unPIC framework for image-to-3D generation. Our choice of intermediate features, CROCS, encodes point-to-point correspondence across views in a canonical reference frame. CROCS allows us to invert the usual “generate then reconstruct” paradigm, by producing 3D directly as part of the view-generation process. With its hierarchical approach and explicit geometric supervision, unPIC outperforms numerous competing methods on novel-view synthesis, multiview-generation consistency, and 3D geometric accuracy. unPIC learns shape priors that generalize well beyond the digital assets it was trained on, paving the way for more geometry-driven novel-view synthesis.
Appendix A Appendix
A.1 CROCS
We first provide a precise specification of CROCS.
Normalization (similar to NOCS)
Let $\Omega\subset\mathbb{R}^{3}$ denote the surface of a 3D object or scene. For any point $\mathbf{p}=(x,y,z)\in\Omega$ in world coordinates, we first normalize its coordinates into NOCS, defined by the unit cube $0,1]^{3}$.
We determine the Axis-Aligned Bounding Box (AABB) of $\Omega$ using the minimum and maximum coordinates, $\mathbf{p}_{\text{min}}$ and $\mathbf{p}_{\text{max}}$. The geometric center of the AABB is $\mathbf{c}_{\Omega}=(\mathbf{p}_{\text{min}}+\mathbf{p}_{\text{max}})/2$. To ensure uniform scaling while preserving the aspect ratio, we calculate the maximum extent $L$ of the bounding box:
$L$ $=\|\mathbf{p}_{\text{max}}-\mathbf{p}_{\text{min}}\|_{\infty}$ (1)
$=\max(x_{max}-x_{min},y_{max}-y_{min},z_{max}-z_{min}).$ (2)
The NOCS coordinates $\mathbf{p}^{\prime}\in[0,1]^{3}$ are then obtained by scaling the object and centering it at $\mathbf{h}=(0.5,0.5,0.5)$:
$\mathbf{p}^{\prime}=\frac{1}{L}(\mathbf{p}-\mathbf{c}_{\Omega})+\mathbf{h}.$ (3)
Camera-Relative Transformation
Let the position of a source camera be defined in spherical coordinates by $(\theta,\phi,r)$, corresponding to the azimuth, elevation, and radial distance, respectively. We assume the camera is always oriented to face the center of the object space.
To achieve camera-relative canonicalization, we rotate the normalized coordinates around the vertical axis (Z-axis) by the source camera’s azimuth $\theta$. This aligns the coordinate system with the viewpoint of the source camera. The transformation is applied around the center of the unit cube $\mathbf{h}$. We define the intermediate rotated coordinates $\mathbf{p}^{\prime\prime}$ as:
$\mathbf{p}^{\prime\prime}=R_{z}(\theta)(\mathbf{p}^{\prime}-\mathbf{h})+\mathbf{h},$ (4)
where $R_{z}(\theta)\in SO(3)$ is the standard 3D rotation matrix around the Z-axis:
$R_{z}(\theta)=\begin{pmatrix}\cos\theta&-\sin\theta&0\\ \sin\theta&\cos\theta&0\\ 0&0&1\end{pmatrix}$. (5)
This transformation rotates the coordinates in the X-Y plane (corresponding to the Red and Green channels in the pointmap) while preserving the Z coordinate (Blue channel). Notably, this definition is deliberately invariant to the camera elevation $\phi$ and distance $r$, a choice further elaborated in Appendix A.2.2.
Rescaling
Since the object was initially scaled to fit an axis-aligned unit cube (Eq. 3), the rotation in Eq. (4) may cause the coordinates $\mathbf{p}^{\prime\prime}$ to exceed the $0,1]^{3}$ bounds (illustrated in Appendix A.2.1). A final rescaling step is necessary to ensure the coordinates remain constrained within the unit cube.
To ensure the normalization is robust and predictable across different azimuth angles, we incorporate the theoretical bounds of the transformation. We define the theoretical range $L_{th}(\theta)$ and the theoretical minimum coordinate $m_{th}(\theta)$ achievable when rotating a unit cube by $\theta$ around its center:
$L_{th}(\theta)$ $=|\cos\theta|+|\sin\theta|,$ (6)
$m_{th}(\theta)$ $=0.5-\frac{1}{2}L_{th}(\theta).$ (7)
Let $\Omega^{\prime\prime}$ be the set of all rotated coordinates for the object. We calculate the observed minimum $m_{obs}$ and maximum $M_{obs}$ values across all points and all dimensions of $\Omega^{\prime\prime}$. The effective minimum shift $m_{eff}$ and the effective range $L_{eff}$ are calculated by clamping the observed values with the theoretical bounds. This ensures that the normalization is robust and consistent with the geometric constraints of the rotation:
$m_{eff}$ $=\max(m_{obs},m_{th}(\theta)),$ (8)
$L_{eff}$ $=\text{clip}(M_{obs}-m_{eff},\epsilon,L_{th}(\theta)),$ (9)
where $\epsilon$ is a small positive constant (e.g., $10^{-6}$) to prevent division by zero.
The final CROCS coordinates $\mathbf{p}_{\text{CROCS}}\in[0,1]^{3}$ are given by the uniformly rescaled coordinates:
$\mathbf{p}_{\text{CROCS}}=\frac{\mathbf{p}^{\prime\prime}-m_{eff}}{L_{eff}}.$ (10)
This procedure guarantees that the object remains tightly bounded by the unit cube in the camera-relative coordinate space.
A.2 Geometric Subtleties
A.2.1 CROCS Rescaling
Recall that we pre-render multiview NOCS images using a fixed reference frame. This reference frame can be labeled using the source-view camera position $(\theta,\phi,r)=(0,0,1)$ that captures the object in its default creator-intended pose at a default camera distance. At training time, we would like to set the source camera to arbitrary locations to ensure a rich training distribution. To change the reference frame from $\theta=0$ to a new source-camera location $\theta^{\prime}$ (we will treat the other parameters in Section A.2.2), we simply rotate the ground-plane axes of all pre-rendered NOCS maps using the SO2 rotation matrix $\cos\theta^{\prime},-\sin\theta^{\prime}],[\sin\theta^{\prime},\cos\theta^{\prime}]]$. The ground-plane axes are the Red and Green channels of the NOCS maps in our case.
Our on-the-fly NOCS-to-CROCS canonicalization comes with a challenge, arising from the fact that the object was scaled to fit the NOCS cube at $\theta=0$. When we rotate the RGB reference cube to follow the primary camera to $\theta^{\prime}$, then the object may no longer be bounded by the rotated cube. See Figure 6 where we visualize this issue for a fixed object scaled to fit the reference cube at $\theta=0$.
In the upper row of Figure 6(a), we get some CROCS values outside the range $[0,1]$, because the object exceeds the boundaries of the $0,1]^{2}$ CROCS reference square. Some values of $\theta^{\prime}$ such as 45, 135, 215, and 305 degrees are affected the most, because they lead to the greatest object overhang. This is not ideal—the scale of CROCS values which the model needs to predict depends on the reference frame. Since the reference frame (i.e., source camera position) is not actually supplied to the model, the model can only learn to predict the variable scale by overfitting. We aim to avoid any dependence on the reference frame. To this end, we introduce a CROCS rescaling operation to eliminate the dependence (Figure 6(a), lower row).
Taking a concrete example, Figure 6(b) shows a cube-like object from different angles, along with the pre-rendered NOCS images (second row), and canonicalized CROCS maps without and with rescaling (third and fourth rows). Here are two observations from this figure: 1) The front-facing edge of the cube, visible at $\theta^{\prime}=45$ as yellow in the second row, disappears in the third row because the CROCS values exceed the $[0,1]$ range (and the color map is defined on the same range). If we rescale the values in the third row, we see the edge reappears in the fourth row. 2) The maximum overhang (in the third row) occurs at $\theta^{\prime}=45$, where the range of observed $y$ values expands to [0.2,1.2][-0.2,1.2]. The length of this expanded interval is nearly $\sqrt{2}\approx 1.41$, which can also be deduced geometrically—it equals the length of the diagonal of the fixed square in Figure 6(a).
Geometrically, we expect no object overhang at all if a given object is bounded by a sphere of diameter 1 (a tighter bound than the unit cube). We believe a lot of objects fall in this category. For this reason, our model does reasonably well even if we train it to predict CROCS without rescaling. But we observed a noticeable gain in the prior’s performance on CROCS with rescaling. The latter setting removes the reference-frame-dependent scale that the model needs to predict in some cases (objects like a cube).
Note that for a given object, the rescaling needs to be performed across all multiview pointmaps (K=8 target views in our case) simultaneously, since any given view only shows a particular facet of the object. (A single pointmap may entirely miss the variation along its depth component.) Our rescaling is reminiscent of the pairwise pointmap renormalization performed by DUSt3R (see Eq. 3 in [43]). All the results in this work use CROCS with rescaling.
(a) The general case (simplified to 2D). Top row: An object that fits the NOCS reference square exactly at $\theta=0$ is no longer bounded by the CROCS reference square at $\theta^{\prime}\in\{15,30,45\}$. Due to object overhang, some CROCS values lie outside $[0,1]$. The extent of the overhang depends on $\theta^{\prime}$. Bottom row: We rescale CROCS based on the observed range of multiview values for any $\theta^{\prime}$. This ensures the object is once again bounded tightly by $0,1]^{2}$.
(b) A concrete example in 3D. We show the minimum and maximum NOCS/CROCS values observed above each pointmap image. At $\theta^{\prime}=45$, we see that the front-facing edge of the cube appears squashed without rescaling, because its X and Y values lie outside the range [0, 1].
Figure 6: CROCS rescaling. We examine the effect of rotating the RGB reference cube used to paint the object’s surface, following the source camera as it moves around a fixed object. $\theta$ (the azimuthal angle) denotes the default camera position, while $\theta^{\prime}$ denotes a new position. We consider the largest possible objects—a square in 2D or cube in 3D.
Figure 7: CROCS when varying the camera elevation angle $\phi$. As in Fig. 3, the large camera denotes the source view; the smaller cameras denote 3 of 7 novel views. The white arrow denotes the normal direction w.r.t. the camera plane (drawn as a white circle connecting the cameras). The colored arrow denotes the normal of the CROCS color cube. Left: at the default angle $\phi=0$, each camera is pointed at a particular face of the RGB color cube, and has a consistent color distribution across examples. Middle: when the source camera is raised to $\phi^{\prime}$, we could hypothetically rotate the color cube by the same degree to ensure the source camera still points at the same face of the cube. The issue is—tilting the color cube toward the source camera tilts it away from the opposite camera, and also rotates the colors seen by the remaining target cameras (by $\phi^{\prime}$ and $\phi^{\prime}$ in their respective image planes). The color shifts experienced by the four cameras are clearly inconsistent. Right: If we choose not to tilt the cube through $\phi^{\prime}$, all cameras undergo a small and consistent shift in the distribution of colors toward the up-axis color (blue). If the model can infer $\phi^{\prime}$, it can also predict the shift. We follow the Right approach.
A.2.2 No Canonicalization for Camera Elevation or Distance
In Section A.2.1 we discussed how to canonicalize NOCS when switching the source camera from $\theta$ to $\theta^{\prime}$. Here, we discuss the remaining camera extrinsic parameters. Since all cameras are oriented to face the center of the object, we do not need to consider their rotations. In fact, we do not need to canonicalize for changes in the camera distance $r$. It only serves as a zoom parameter to augment the training data.
The only remaining parameter is the elevation angle $\phi$, which determines the source-camera height. Recall that the target cameras are also placed at the same camera height. In fact, $\phi$ determines the difficulty of the multiview prediction problem. For instance, at $\phi=90$ (a straight-down view of the object), the source and target camera locations converge to the same point—the apex of the hemisphere. In this case, the prediction task is reduced to rotating the object in the image plane.
Say we raise/lower the source camera from $\phi$ to $\phi^{\prime}$ (Figure 7). This would reveal more/less of the object’s upper/lower surface, and hence more/less of the NOCS color representing the up-axis (Blue in our case). We could potentially correct for this by tilting the reference cube so the source camera is pointed at the same cube face as before at $\phi$. This would help maintain the same color bias for the source camera. However, the target cameras would no longer point at the same cube faces respectively—rather, their distribution of colors would shift (toward the up-axis/away from it in the case of the opposite target camera) or rotate (in the image plane for other target cameras). Consequently, the distribution shifts in the CROCS colors would be inconsistent across target cameras.
Based on this geometrical observation, we choose not to canonicalize for changes in the camera elevation $\phi^{\prime}$. This has the effect of leaving it to the model to infer the camera elevation from the source image (implicitly in contrast to One-2-3-45 [26]). It differs from our treatment of $\theta^{\prime}$, which leaves the model oblivious to the azimuthal rotation of the cameras.
Our decision not to canonicalize for $\phi$ is forced by the design choice of our target camera poses, which follow the camera height of the source camera. An alternative choice would be to tilt the camera plane and the CROCS cube together, leaving the cameras at different heights in Figure 7-Middle [^1] This alternative is equivalent to using a stationary multiview camera rig, but orienting the object randomly before taking pictures, as in CAT3D [10], likely LRM [15], and datasets like YCB [2]. Changing the object’s orientation decouples the object’s up-axis from the world’s up-axis. It can leave the object in unnatural, physically unstable poses.. Our choice of target poses is based on how humans perceive and reason about objects. We tend to see side or top-down views of level objects. On mental rotation tasks (e.g., when asked what an object looks like from “behind”), we tend to imagine a rotation of the object either in the image plane, or in depth about the world’s vertical axis [36, 39]. We find it harder to imagine a potentially bottom-facing view. In addition to aligning with human perception, modeling object rotations at typical views/poses (e.g., top-down or side angles) could also be more useful for downstream applications.
Appendix B Training and Evaluation Details
B.1 Dataset
We use the following sets of assets from Objaverse: 1) the LVIS subset expanded to 83k examples from the same categories. We use object type labels predicted with high confidence from [17] to expand the LVIS subset. 2) The KIUI subset  [^2] https://github.com/ashawkey/objaverse_filter comprising 101k additional assets. 3) From the Objaverse-XL alignment subset [4], we used a combination of GLB, OBJ, and FBX assets after filtering out certain (a) terrains and HDRI environment maps, (b) layouts and rooms, and (c) textureless FBX assets that rendered as pink. In total, we used 620k assets (Objaverse and Objaverse-XL) for training. In addition, we held out 50k assets randomly sampled from Objaverse-XL after the filtering step.
To render NOCS images, we adapted a script from BlenderProc [^3] https://github.com/DLR-RM/BlenderProc/ [5] which creates a special NOCS material for the surface of a given object. We also use their Blender settings (the CYCLES engine with 1 diffuse bounce, 0 glossy bounces, and 0 ambient occlusion bounces) for rendering NOCS. We export them in the EXR format to ensure linear-light values, preserving the continuity of NOCS values on the object’s surface.
B.2 Model Hyperparameters
[…]
We train the prior and decoder modules with a conditioning dropout probability of 0.08 on the source image. We do not use any dropout on the CROCS images fed to the decoder to ensure it adheres to the geometry. At test time, we use a fixed classifier-free guidance weight of 2.0 on the source image for both modules. We use an exponential moving average of the model parameters for evaluation, although this is not essential.
The unPIC base model has a total of 1.1B parameters across the prior (470M), decoder (470M), and two VAEs (84M each). It takes about 1min 13s to run all model components, with classifier-free guidance over 1000 denoising steps, on an A100 GPU. Note that we did not attempt to improve the runtime, as that was not the core aim of this work. Techniques such as step distillation, or recent advances in Optimal Flow Matching [22] which facilitate straight trajectories for fast denoising, would be readily applicable to unPIC.
B.3 Choice of K=8 Views
To motivate why $K=8$ views are sufficient to produce a good high-level 3D reconstruction, we take a larger number of ground-truth views ($K=24$), and look at the pairwise similarity between them. This can be computed using CLIP embeddings (like our CLIP-MV metric). See Figure 8 for a heatmap of pairwise similarity scores.
The black crosses that occur in the heatmap show the frequency at which rotated views become dissimilar to each other. We see  6 black crosses on either side of the diagonal of the matrix. This number estimates how many views are sufficient to summarize the full 360-degree rotation. Hence, by generating 7 novel views ($K=8$) for any input image, we can capture 3D structure at sufficient granularity for reconstruction.
Note that this analysis does not take into account top and bottom views of the object.
Figure 8: Pairwise cosine similarities of CLIP embeddings for rotations of a given object, averaged across a subset of ObjaverseXL. The x- and y-axes both show 24 views across a 360-degree rotation.
B.4 U-Net Architecture
The denoising network follows a standard design with a downsampling encoder path, a bottleneck middle path, and an upsampling decoder path with skip connections. The exact architecture is as follows, while the hyperparameters are summarized in Table 6:
  • Input: A noisy latent superimage $X$ and a time-step embedding $t_{emb}$ derived from the diffusion time $t$. Any conditioning superimages (e.g., CROCS latents, $Z^{g}$) are concatenated with $X$ along the channel axis.
[…]
  • Middle Stack: This path consists of three Resnet Blocks with Multi-Head Attention Blocks in between them to process the lowest-resolution feature map.
  • Upsampling Stack: This path mirrors the downsampling path. At each level, the feature map is upsampled, concatenated with the corresponding skip connection from the downsampling path, and passed through three Resnet Blocks. A Multi-Head Attention Block is also applied at specified resolutions.
  • Output: The final feature map is processed through a normalization layer, a Swish activation, and a final $\times3$ convolution to produce the denoised latent superimage of the same shape as the input $X$.
Core Components:
  • Resnet Block: A standard residual block containing two $\times3$ convolutional layers, Group Normalization, and Swish activations. The time-step embedding $t_{emb}$ is projected and used to perform a feature-wise affine transformation (scale and shift) on the feature map after the first normalization layer.
  • Multi-Head Attention Block: A standard multi-head self-attention mechanism applied over the spatial dimensions of the feature map. This allows every location in the feature map to integrate information from all other locations, enabling global communication across the different views tiled in the superimage.
Table 6: U-Net Hyperparameters
Parameter Value
embed_dim 256
channel_mult (1, 2, 2, 3, 3)
num_res_blocks 3
per_head_channels 64
unet_dropout 0.0
norm_type ’Group Norm’
B.5 VAE Fine-Tuning Metrics
We fine-tuned the Stable Diffusion 1.4 VAE, on CROCS and RGB images separately, using a combination of losses: Reconstruction, LPIPS, KL-Divergence, and the GAN loss. See Table 7 showing the improvement in each VAE’s metrics after fine-tuning. The CROCS VAE benefits significantly, and outperforms the RGB VAE in terms of accuracy.
Table 7: VAE metrics on ObjaverseXL before and after fine-tuning on (a) CROCS pointmaps and (b) RGB images.
RGB VAE CROCS VAE
Before Fine-Tuning PSNR 37 38
FID 62 16
KL 28005 20823
After Fine-Tuning PSNR 38 51
FID 16 2
KL 87496 35936
B.6 Assessing Multiview Consistency using MET3R
To assess multiview consistency, we attempted to run MET3R, which is a pose-free way to evaluate the consistency of generated novel views or frames of a generated video. Using ground-truth images, we found that even a 45-degree rotation of an object is too significant for MET3R. The metric produced values (see Fig 9) exceeding the expected value of the metric on random noise (around 0.3). We ran with the standard settings, using the MAST3R backbone, DINO16 features, FeatUp upsampling, and the cosine distance.
Figure 9: MET3R as a metric for multiview consistency produced values in excess of those expected from random noise, even on ground-truth images. The 45-degree rotation in each case is too great to facilitate a dense 3D reconstruction.
B.7 Baselines
For single-image-to-multiview synthesis, we compare unPIC with the following baselines:
  1. 1.
    CAT3D (512x512): a SoTA-quality, geometry-free diffusion model trained using masked modeling of views on diverse datasets. While the original model is closed source, we received a Colab and checkpoint from the authors. The model uses a coarse view-sampling strategy to generate 7 anchor views given 1 source image. Subsequently, it conditions on the anchor views to generate a finer set of views using the same model. We focus on the coarse stage for our evaluation.
  2. 2.
    EscherNet (256x256): a geometry-free diffusion model equipped with a specialized camera encoding (CaPE), allowing a flexible number of input images and target views. It uses the pretrained StableDiffusion v1.5 model as its latent diffusion backbone.
  3. 3.
    Free3D (256x256): a geometry-free multiview diffusion model that uses Plücker rays to encode target camera poses and modulate the diffusion latents. It is trained on Objaverse only.
  4. 4.
    InstantMesh (512x512): a geometry-aware, feed-forward framework that generates a 3D mesh from a single image. We use the Large variant of the model. InstantMesh first employs a multi-view diffusion model, Zero123++, to produce a set of 3D-consistent views from the input. These views are then processed by a transformer-based sparse-view Large Reconstruction Model (LRM) that directly outputs a mesh. The model was trained with direct geometric supervision (depths and normals) on the mesh representation using a differentiable iso-surface extraction module (FlexiCubes). We use the frames rendered from the 3D mesh for our evaluation.
  5. 5.
    One-2-3-45 (256x256): a SoTA version of Zero-1-to-3 [27]. Its diffusion-based image prior takes camera rotation and translation (R and T) parameters to transform a given image. Each novel view is sampled individually, leaving the outputs prone to multiview inconsistencies. We focus on evaluating the geometry-free stage, feeding R and T parameters to match our target poses, but using the model’s own estimation of the camera height.
  6. 6.
    OpenLRM (288x288): a geometry-aware but deterministic method based on the original LRM [15]. It uses a Transformer to convert DINO features into implicit triplane representations of 3D shape. The triplane tokens are spatially indexed/queried to parameterize a NeRF model of the scene, mitigating view consistency issues and allowing image-generation at arbitrary camera poses.
Appendix C Additional Results
Task 1: Novel-view synthesis. We provide additional multiview comparisons between unPIC and the baselines in Figure 10. We show five examples at eight target views each.
Task 2: 3D reconstruction. We visualize our point clouds and compare them with two baselines in Figure 15. We also show meshes obtained by applying a surface-reconstruction algorithm (Screened Poisson [18]) on our point clouds. We show 40 examples at one novel view each.
In-the-wild generalization. We apply unPIC to arbitrary real-world images in Figure 23. We also cover failure cases in Figure 24.
Figure 10: Novel views from baselines versus our method (figure continues on next page). We show an example from our ObjaverseXL test set.
Figure 11: Novel views from baselines versus our method (figure continues on next page). We show an example from our ObjaverseXL test set.
Figure 12: Novel views from baselines versus our method (figure continues on next page). We show an example from our Amazon Berkeley Objects test set.
Figure 13: Novel views from baselines versus our method (figure continues on next page). We show an example from our Digital Twin Catalog test set.
Figure 14: Novel views from baselines versus our method (continued). We show an example from our Digital Twin Catalog test set.
(a) Example 1
(b) Example 2
(c) Example 3
Figure 15: 3D reconstruction results from unPIC vs baseline methods (Part 1 of 8). We show a $195^{\circ}$ rotation of the object relative to the input image. The bottom row (unPIC’s reconstructed mesh) is visualized with transparency to show the full 3D shape. (Figure continues on next page.)
(a) Example 4
(b) Example 5
(c) Example 6
Figure 16: 3D reconstruction results from unPIC vs baseline methods (Part 2 of 8). We show a $195^{\circ}$ rotation of the object relative to the input image. The bottom row (unPIC’s reconstructed mesh) is visualized with transparency to show the full 3D shape. (Figure continues on next page.)
(a) Example 7
(b) Example 8
(c) Example 9
Figure 17: 3D reconstruction results from unPIC vs baseline methods (Part 3 of 8). We show a $195^{\circ}$ rotation of the object relative to the input image. The bottom row (unPIC’s reconstructed mesh) is visualized with transparency to show the full 3D shape. (Figure continues on next page.)
(a) Example 10
(b) Example 11
(c) Example 12
Figure 18: 3D reconstruction results from unPIC vs baseline methods (Part 4 of 8). We show a $195^{\circ}$ rotation of the object relative to the input image. The bottom row (unPIC’s reconstructed mesh) is visualized with transparency to show the full 3D shape. (Figure continues on next page.)
(a) Example 13
(b) Example 14
(c) Example 15
Figure 19: 3D reconstruction results from unPIC vs baseline methods (Part 5 of 8). We show a $195^{\circ}$ rotation of the object relative to the input image. The bottom row (unPIC’s reconstructed mesh) is visualized with transparency to show the full 3D shape. (Figure continues on next page.)
(a) Example 16
(b) Example 17
(c) Example 18
Figure 20: 3D reconstruction results from unPIC vs baseline methods (Part 6 of 8). We show a $195^{\circ}$ rotation of the object relative to the input image. The bottom row (unPIC’s reconstructed mesh) is visualized with transparency to show the full 3D shape. (Figure continues on next page.)
(a) Example 19
(b) Example 20
(c) Example 21
Figure 21: 3D reconstruction results from unPIC vs baseline methods (Part 7 of 8). We show a $195^{\circ}$ rotation of the object relative to the input image. The bottom row (unPIC’s reconstructed mesh) is visualized with transparency to show the full 3D shape. (Figure continues on next page.)
(a) Example 22
(b) Example 23
(c) Example 24
Figure 22: 3D reconstruction results from unPIC vs baseline methods (Part 8 of 8). We show a $195^{\circ}$ rotation of the object relative to the input image. The bottom row (unPIC’s reconstructed mesh) is visualized with transparency to show the full 3D shape. (Figure continues on next page.)
Figure 23: unPIC on in-the-wild images. Each pair of rows shows: the input image in the top-left box, predicted CROCS in the middle (2x4 grid), and predicted novel-view images to the right (2x4).
Figure 24: Failure cases on in-the-wild images. Each pair of rows shows: the input image in the top-left box, predicted CROCS in the middle (2x4 grid), and predicted novel-view images to the right (2x4). In the first example, unPIC rotates the person in the image plane, likely due to interpreting the source image as an overhead view. In the second example, unPIC predicts plausible poses, but fails to texture the images correctly, likely due to the out-of-distribution case of multiple humans in the scene.